# Transfer Learning of Surrogate Models via Domain Affine Transformation Across Synthetic and Real-World Benchmarks

### v1

@@ Abstract
Surrogate models are frequently employed as efficient substitutes for the costly execution of real-world processes. However, constructing a high-quality surrogate model often demands extensive data acquisition. A solution to this issue is to transfer pre-trained surrogate models for new tasks, provided that certain invariances exist between tasks. This study focuses on transferring non-differentiable surrogate models (e.g., random forest) from a source function to a target function, where we assume their domains are related by an unknown affine transformation, using only a limited amount of transfer data points evaluated on the target. Previous research attempts to tackle this challenge for differentiable models, e.g., Gaussian process regression, which minimizes the empirical loss on the transfer data by tuning the affine transformations. In this paper, we extend the previous work to the random forest model and assess its effectiveness on a widely-used artificial problem set - Black-Box Optimization Benchmark (BBOB) testbed, and on four real-world transfer learning problems. The results highlight the significant practical advantages of the proposed method, particularly in reducing both the data requirements and computational costs of training surrogate models for complex real-world scenarios.
@@
transfer learning, random forest, Gaussian process, real-world applications
@@ I Introduction
Surrogate modelling [1, 2, 3, 4] is commonly utilized across various fields of computer science and engineering, particularly for complex simulations or physical experiments where data acquisition is expensive. Common learning algorithms for building surrogate models include artificial neural networks [5], support vector machine [6, 7], random forest regression [8, 9] and Gaussian process regression [10, 11, 12]. These learning algorithms often have a high sample complexity, especially when the dimension of the independent variable becomes large. As such, when surrogate-modeling a new regression task, one might wish to exploit surrogate models trained on previous tasks, given similarities and symmetries between tasks.
@@
This paper is organized as follows. In Sec. II, we summarize some key related works. In Sec. III, we delineate our extension of the affine transformation method to non-differentiable functions. Experimental setup and both synthetic and real-world transfer learning tasks are explained in Sec. IV, followed by a detailed discussion of the results in Sec. V. Finally, we conclude the empirical findings and point out the future direction in Sec. VI.
@@ III Affine Transfer Learning for Non-differentiable Models
We propose to use the Covariance matrix adaptation evolution strategy (CMA-ES) [20, 21, 22] to tune the affine parameters. CMA-ES can be applied directly to the search space of the translation parameter $\mathbf{v}$, which is Euclidean. However, special treatment is needed for $\mathbf{W}$, which lives in a smooth manifold $\operatorname{SO}\left(d\right)$. To solve this issue, we consider the Lie group representation $\mathfrak{so}(d)=\{\mathbf{A}\in\mathbb{R}^{d\times d}\colon\mathbf{A}^{\top}=-\mathbf{A}\}$, which is a flat space (with dimension $d(d-1)/2$), and optimize this representation with CMA-ES. A rotation matrix $\mathbf{W}$ can be recovered its representation $\mathbf{A}$ with the exponential map, i.e., $\mathbf{W}=\operatorname{Exp}(\mathbf{A})$. For each search point $\mathbf{z}\in\mathbb{R}^{d(d-1)/2}$, we have to transform it into a $d\times d$ antisymmetric matrix to preserve the structure of $\mathfrak{so}(d)$: the components in $\mathbf{z}$ are sequentially assigned to the upper triangular of $\mathbf{A}$ (diagonal entries are zero) row by row. The lower triangular entries are then filled by the negative value of the transposition of the upper triangular. We summarize this Riemannian version of CMA-ES in Alg. 1.
@@ 1 Procedure: $\mathcal{L}$, $\lambda$, $\sigma$);
3 Sample $\mathbf{z}$ u.a.r. in $\mathbb{R}^{d(d-1)/2}$ and $\mathbf{v}$ u.a.r. in $\mathbb{R}^{d}$;
@@ IV-B1 Porkchop Plot Benchmarks in Interplanetary Trajectory Optimization
The energy required to travel between Earth and another planet in the solar system is a function of the relative positions and velocities of the planets, as well as the desired trip duration, i.e., of both the departure and arrival times. Calculating this energy for specific departure and arrival dates involves solving Lambert’s problem, a differential equation that determines the orbital trajectory [28]. The periodic motion of planets generates recurring patterns in the contour lines of this function around local minima. These contours resemble the shape of a porkchop slice, inspiring the term “porkchop plots” [29, 30]. The analysis of porkchop plots is useful in the design of interplanetary trajectories. For benchmarking, we consider as the source function the energy required for an Earth-to-Mars mission within a specific launch window and within a maximum travel time of 14 years and the target function corresponds to a different launch window. In this way, we aim to transfer insights gained from the planning of an interplanetary mission during a specific launch window to plan similar missions for future windows. Besides the Earth-to-Mars mission, we also validate our transfer learning approach on Earth-to-Venus (a maximum travel time of 12 years) and Mercury-to-Earth (a maximum travel time of 6 years) missions.
@@ IV-B3 Real-World Optimization Benchmark in Vehicle Dynamics
Figure 2: The comparison evaluates the performance of Random forest regression models obtained through transfer learning against those trained from scratch on the transfer dataset. This analysis considers various transfer data sample sizes and problem dimensions. Each cell in the figure shows the percentage difference in average SMAPE (%) between the two approaches for specific BBOB functions, dimensions, and sample sizes. Positive values (marked in red) indicate that the transferred model achieves superior accuracy by yielding a lower SMAPE compared to the model trained from scratch.
@@ IV-C Surrogate model
In our experiments, we evaluate the performance of RFR using the SMAPE [34], chosen for its scale-neutral properties that enable fair comparisons across datasets with different scales. Transfer learning experiments are repeated ten times, with randomly generated affine transformations applied to each BBOB function to validate our method. Similarly, real-world applications are also tested ten times to account for variability and ensure robust and consistent results.
@@ IV-D Implementation details
Experiments are conducted using Python 3.9, with BBOB functions evaluated through the IOHexperimenter framework [35]. RFR is implemented using the scikit-learn library [36] with default hyperparameters when validating on the BBOB functions, while CMA-ES is implemented with the pycma library [37]. For the vehicle dynamics task. hyperparameter tuning is performed with SMAC3 [38] for the original surrogates, optimizing parameters such as the number of trees ([100,1 500]1001500[100,1\,500][ 100 , 1 500 ]), maximum tree depth ($[10,60]$), minimum samples for node splitting ($[2,20]$), leaf node samples ($[1,10]$), and features fractions ($[0.1,1.0]$). Default settings with 500 trees are used for other real-world tasks. To address the skewed distribution of relative function values (differences from the global optimum) in BBOB functions, a $\ log$ transformation is applied before fitting the surrogate model, a step also used in real-world applications. For the Kinematics of a Robot Arm, we follow the pre-processing steps outlined in [39]. The poliastro library [40] is used for generating Porkchop Plot benchmarks, with data generation details for other applications available in their respective publications.
@@ IV-E Reproducibility
The complete experimental setup and algorithm implementation are available in our Zenodo repository [41]. The repository also includes all necessary files and datasets to ensure reproducibility. Additionally, it provides a copy of the supplementary materials for convenient access.
@@ V Experimental Results
Figure 3: The comparison evaluates the performance of Random forest regression models obtained through transfer learning against those trained from scratch on the Porkchop Plot Benchmarks in Interplanetary Trajectory Optimization, focusing on the Earth-to-Mars mission. This analysis examines various transfer data sample sizes. Each cell in the figure shows the percentage difference in average SMAPE (%) between the two approaches for specific transfer settings and sample sizes. Positive values indicate that the transferred model achieves superior accuracy by yielding a lower SMAPE compared to the model trained from scratch.
@@ V-A Transferring RFR on BBOB
Fig. 2 highlights the advantages of our approach by presenting the average SMAPE differences (percentage) between the RFR trained from scratch and the transferred RFR. For the majority of functions, the transferred model consistently provides a better fit, as indicated by positive differences. Notably, the performance gap narrows with increasing sample size, indicating that the benefits of transfer learning diminish as more data becomes available, allowing the model trained from scratch to catch up. Compared to the transferred GPR results reported in the prior study [13], the RFR exhibits more stable performance, with no extreme outliers observed in the 10-dimensional (10-D) case using 50 transfer samples. This demonstrates that the RFR can still deliver reasonable results even when data is sparse in high-dimensional scenarios. However, achieving effective transfer for certain functions remains challenging. For instance, across various dimensions and transfer data sizes, the transferred RFR consistently struggles within function F24, where the model trained from scratch consistently outperforms it. Similarly, functions F21–F23, especially in high-dimensional scenarios, present substantial challenges for the transferred model, limiting its ability to deliver optimal performance.
@@
For a detailed analysis of the 5-D case, TABLE II (refer to the supplementary material [41]) reports the SMAPE values (mean and standard deviations across ten runs) for three models: the original RFR, the transferred model, and the one trained from scratch. Similar results for the 2-D and 10-D are included in the supplementary material [41]. The table reports outcomes for two different sizes of the transfer data: $\lvert\mathcal{T}\rvert\in\{50,50d\}$. Statistical comparisons between the models are conducted using the Kruskal-Wallis test, followed by Dunn’s posthoc analysis, with a significance level of 5%. To indicate statistical significance, we highlight the following: if the transferred model demonstrates superior performance compared to the original model, it is underlined. Furthermore, when comparing the transferred model with the model trained from scratch, the better-performing model is highlighted in bold.
In 5-D functions, the transferred RFR generally outperforms the original model and the model trained from scratch at transfer data sizes of 50 and 250, with rare exceptions. An exception arises with F23, where the original model outperforms both the transferred models and the one trained from scratch. Specifically, when the transfer dataset is 50, the original RFR demonstrates significantly better performance than the transferred RFR. Further analysis reveals that the original RFR trained on the source function exhibits considerable underfitting - a behavior also observed with the original GPR model [13]. In such cases, an affine transformation is unlikely to improve the model’s performance on the target function effectively. These findings suggest that the proposed transfer learning approach is most effective when the original surrogate model already demonstrates a reasonably good fit to the source function.
@@
Figure 7 (see the supplementary material [41]) provides a detailed visualization of how model performance varies with the size of the transfer dataset for the 5-D case. The SMAPE trends indicate that, for most functions, the transferred model outperforms the model trained from scratch up to a transfer dataset size of 250 points. This trend supports the expected suitability of our transfer learning method for scenarios with relatively small sample sizes. However, there are notable exceptions. For function F24, the model trained from scratch surpasses the transferred model with as few as 40 samples. Similarly, for F21, the model trained from scratch consistently outperforms the transferred model regardless of dataset size. Additionally, in the cases of F16 and F23, the original RFR demonstrates superior performance, consistently exceeding both the transferred model and the model trained from scratch across all transfer dataset sizes, ranging from 10 to 250 samples.
Next, we explore the impact of the function’s domain dimensionality on the performance and efficacy of the proposed transfer learning approach. In 2-D functions (see Fig. 6 and Table I in the supplementary material [41]), the transferred model demonstrates superior performance over the model trained from scratch for most functions at smaller sample sizes (e.g., 50), with F23 and F24 being notable exceptions. This advantage largely persists as the number of transfer samples increases, although F19 and F24 emerge as additional exceptions. Furthermore, the transferred model consistently outperforms the original model across all tested sample sizes, except for function F23. In 10-D functions (refer to the supplementary material [41]), an intriguing trend emerges as the transfer dataset size increases. With smaller datasets (e.g., 40 samples), the transferred RFR significantly outperforms the model trained from scratch, highlighting a clear performance advantage. However, as the dataset size grows (e.g., 500 samples), this gap gradually diminishes, reflecting the scratch-trained model’s ability to catch up with more data. Despite this, transfer learning proves less effective for certain highly multi-modal functions, such as F16 and F21 - F24, where the inherent complexity poses persistent challenges.
@@ V-B Transferring RFR on Real-world applications
Figure 4: The comparison evaluates the performance of Random forest regression models obtained through transfer learning against those trained from scratch on the Kinematics of a Robot Arm. This analysis examines various transfer data sample sizes. Each cell in the figure shows the percentage difference in average SMAPE (%) between the two approaches for specific transfer settings and sample sizes. Positive values indicate that the transferred model achieves superior accuracy by yielding a lower SMAPE compared to the model trained from scratch.
@@ V-B1 Porkchop Plot Benchmarks in Interplanetary Trajectory Optimization
Fig. 3 illustrates the percentage difference in average SMAPE between RFR models trained from scratch and those adapted through transfer learning, evaluated across various transfer scenarios and transfer dataset sizes. The result reveals that, in most cases, the transferred model outperforms the model trained from scratch, particularly when the transfer dataset contains fewer than 100 samples. This indicates that transfer learning is especially advantageous in data-scarce settings. However, as the size of the transfer dataset increases, the performance gap diminishes, with the models trained from scratch showing greater improvements. A more detailed analysis of the relation between performance and transfer dataset size is provided in Fig. 9 (see the supplementary material [41]). In most scenarios, the transferred RFR consistently outperforms the original RFR, and the model is trained from scratch when only a small transfer dataset is available. However, as the dataset grows, the from-scratch model outperforms the transferred model.
The original RFR model rarely surpasses the transferred or scratch-trained models when the transfer dataset is extremely limited, such as with just 10 samples, though occasional exceptions are observed. In such exceptions, the transferred model may overfit the limited data during optimization, leading to poor generalization to the broader target domain. Overfitting can be identified when the transferred model performs better than the original RFR on the transfer samples, as it is explicitly optimized for them. These findings highlight that while the transfer learning approach is generally robust, its effectiveness can be challenged in scenarios with severely restricted data availability. Similar performance patterns are observed in other experimental scenarios, such as Earth-to-Venus and Mercury-to-Earth transfers, as shown in Fig. 10, 11, 12, and 13, which are included in the supplementary material [41].

### v3

@@ Abstract
Surrogate models are frequently employed as efficient substitutes for the costly execution of real-world processes. However, constructing a high-quality surrogate model often demands extensive data acquisition. A solution to this issue is to transfer pre-trained surrogate models for new tasks, provided that certain invariances exist between tasks. This study focuses on transferring non-differentiable surrogate models (e.g., random forests) from a source function to a target function, where we assume their domains are related by an unknown affine transformation, using only a limited amount of transfer data points evaluated on the target. Previous research attempts to tackle this challenge for differentiable models, e.g., Gaussian process regression, which minimizes the empirical loss on the transfer data by tuning the affine transformations. In this paper, we extend the previous work to the random forest and assess its effectiveness on a widely-used artificial problem set - Black-Box Optimization Benchmark (BBOB) testbed, and on four real-world transfer learning problems. The results highlight the significant practical advantages of the proposed method, particularly in reducing both the data requirements and computational costs of training surrogate models for complex real-world scenarios.
@@
transfer learning, random forest, real-world applications
publicationid: pubid: 979-8-3315-3431-8/25/\$31.00 ©2025 IEEE
@@ I Introduction
Surrogate modeling [1, 2, 3, 4] is commonly utilized across various fields of computer science and engineering, particularly for complex simulations or physical experiments where data acquisition is expensive. Common learning algorithms for building surrogate models include artificial neural networks [5], support vector machine [6, 7], random forests [8, 9] and Gaussian process regression [10, 11, 12]. These learning algorithms often have a high sample complexity, especially when the dimension of the independent variable becomes large. As such, when surrogate-modeling a new regression task, one might wish to exploit surrogate models trained on previous tasks, given similarities and symmetries between tasks.
@@ III Affine Transfer Learning for Non-differentiable Models
We propose to use Covariance matrix adaptation evolution strategy (CMA-ES) [20, 21, 22] to tune the affine parameters. CMA-ES can be applied directly to the search space of the translation parameter $\mathbf{v}$, which is Euclidean. However, special treatment is needed for $\mathbf{W}$, which lives in a smooth manifold $\operatorname{SO}\left(d\right)$. To solve this issue, we consider the Lie group representation $\mathfrak{so}(d)=\{\mathbf{A}\in\mathbb{R}^{d\times d}\colon\mathbf{A}^{\top}=-\mathbf{A}\}$, which is a flat space (with dimension $d(d-1)/2$), and optimize this representation with CMA-ES. A rotation matrix $\mathbf{W}$ can be recovered its representation $\mathbf{A}$ with the exponential map, i.e., $\mathbf{W}=\operatorname{Exp}(\mathbf{A})$. For each search point $\mathbf{z}\in\mathbb{R}^{d(d-1)/2}$, we have to transform it into a $d\times d$ antisymmetric matrix to preserve the structure of $\mathfrak{so}(d)$: the components in $\mathbf{z}$ are sequentially assigned to the upper triangular of $\mathbf{A}$ (diagonal entries are zero) row by row. The lower triangular entries are then filled by the negative value of the transposition of the upper triangular. We summarize this Riemannian version of CMA-ES in Alg. 1.
@@ 1 Procedure: $\mathcal{L}$, $\lambda$, $\sigma$);
3 Sample $\mathbf{z}$ uniformly at random in $\mathbb{R}^{d(d-1)/2}$ and $\mathbf{v}$ uniformly at random in $\mathbb{R}^{d}$;
@@ IV-B1 Porkchop Plot Benchmarks in Interplanetary Trajectory Optimization
The energy required to travel between Earth and another planet in the solar system is a function of the relative positions and velocities of the planets, as well as the desired trip duration, i.e., of both the departure and arrival times. Calculating this energy involves solving Lambert’s problem, a differential equation that determines the orbital trajectory [28]. The periodic motion of planets generates recurring patterns in the contour lines of this function around local minima. These contours resemble the shape of a porkchop slice, inspiring the term “porkchop plots” [29, 30]. The analysis of porkchop plots is useful in the design of interplanetary trajectories. For benchmarking, we consider as the source function the energy required for an Earth-to-Mars mission within a specific launch window and within a maximum travel time of 14 years and the target function corresponds to a different launch window. In this way, we aim to transfer insights gained from the planning of an interplanetary mission during a specific launch window to plan similar missions for future windows. Besides the Earth-to-Mars mission, we also validate our approach on Earth-to-Venus (a maximum travel time of 12 years) and Mercury-to-Earth (a maximum travel time of 6 years) missions.
@@ IV-B3 Real-World Optimization Benchmark in Vehicle Dynamics
Figure 2: The comparison evaluates Random forest regression models obtained through transfer learning against those trained from scratch on the transfer dataset, across varying sample sizes and problem dimensions. Each cell in the figure shows the percentage difference in average SMAPE (%) between the two approaches for specific BBOB functions. Positive values (marked in red) indicate better accuracy from transfer learning (i.e., lower SMAPE).
@@ IV-C Surrogate model
We evaluate RFR using SMAPE [34] for its scale-invariant properties, allowing fair comparisons across datasets. Transfer learning experiments are repeated ten times with random affine transformations applied to each BBOB function. Similarly, real-world applications are also tested ten times to account for variability and ensure robust and consistent results.
@@ IV-D Implementation details
Experiments are conducted using Python 3.9 using BBOB functions via the IOHexperimenter framework [35]. RFR is implemented with scikit-learn library [36] using default hyperparameters when validating on BBOB functions. CMA-ES is implemented with pycma [37], where the parameter vector encodes rotation and translation. Rotation is initialized by sampling a Gaussian antisymmetric matrix, applying the exponential map, and extracting the upper-triangular elements. Translation is randomly sampled from $[-0.5,0.5]$ per dimension. Bounds are set to $[-\pi,\pi]$ for rotation and $[-1.5,1.5]$ for translation, with the initial step-size set to one-fifth of the average range. All other settings follow pycma defaults, using BIPOP-CMA-ES with three restarts for robustness.
For the vehicle dynamics task, RFR hyperparameters are tuned using SMAC3 [38], optimizing tree count ([100,1 500]1001500[100,1\,500][ 100 , 1 500 ]), depth ($[10,60]$), split/leaf samples ($[2,20]$, $[1,10]$), and features fractions ($[0.1,1.0]$). Other real-world tasks use default settings with 500 trees. To address the skewed distribution of relative function values (differences from the global optimum) in BBOB functions, a $\ log$ transformation is applied before model fitting. For the Robot Arm Kinematics task, we follow the pre-processing [39]. Porkchop Plot benchmarks are generated using poliastro [40], with other data sourced from their respective references.
@@ IV-E Reproducibility
The complete experimental setup, implementation, and supplementary materials are available in our Zenodo repository [41] for full reproducibility.
@@ V Experimental Results
Figure 3: The comparison evaluates Random forest regression models obtained through transfer learning against those trained from scratch on Porkchop Plot Benchmarks for Earth-to-Mars trajectory optimization. It examines different transfer data sample sizes, with each figure cell showing the percentage difference in average SMAPE (%) between the two approaches for specific transfer settings. Positive values indicate better accuracy from transfer learning (i.e. lower SMAPE).
@@ V-A Transferring RFR on BBOB
Fig. 2 shows the average SMAPE differences (%) between RFR trained from scratch and the transferred RFR. In most cases, the transferred model achieves better fits, as indicated by positive differences. Notably, the performance gap narrows with increasing sample size, indicating that the benefits of transfer learning diminish as more data becomes available, allowing the model trained from scratch to catch up. Compared to the transferred GPR results reported in the prior study [13], the RFR exhibits more stable performance, with no extreme outliers observed in the 10-dimensional (10-D) case using 50 transfer samples. This demonstrates that the RFR can still deliver reasonable results even when data is sparse in high-dimensional scenarios. However, achieving effective transfer for certain functions remains challenging. For instance, across various dimensions and transfer data sizes, the transferred RFR consistently struggles within function F24, where the model trained from scratch consistently outperforms it. Similarly, functions F21–F23, especially in high-dimensional scenarios, present substantial challenges for the transferred model, limiting its ability to deliver optimal performance.
@@
For a detailed analysis of the 5-D case, TABLE II (see supplementary material [41]) reports mean and standard deviation of SMAPE across ten runs for three models: original RFR, transferred RFR, and RFR trained from scratch. Results for 2-D and 10-D are also provided in the supplementary material [41]. Two transfer data sizes are evaluated: $\lvert\mathcal{T}\rvert\in\{50,50d\}$. Statistical significance is assessed using the Kruskal-Wallis test followed by Dunn’s posthoc test at the 5% level. In the table, transferred models outperforming the original are underlined; bold highlights indicate the better model between the transferred model with the model trained from scratch.
In 5-D functions, the transferred RFR generally outperforms both the original and scratch-trained models at transfer sizes of 50 and 250, with rare exceptions. Notably, for F23 with 50 transfer samples, the original model significantly outperforms the transferred one. Analysis shows that the original RFR underfits the source function, making the affine transfer ineffective. These results indicate that transfer learning is most beneficial when the original surrogate already fits the source function reasonably well.
@@
Figure 8 (see supplementary material [41]) visualizes how model performance in the 5-D case varies with transfer dataset size. For most functions, the transferred RFR outperforms the model trained from scratch up to 250 samples, confirming its effectiveness in low-data regimes. However, exceptions exist: for F24, the model trained from scratch outperforms the transferred one with as few as 40 samples; for F21, it consistently performs better across all sizes. In F16 and F23, the original RFR outperforms both the transferred model and the model trained from scratch across all dataset sizes (10–250 samples).
Next, we examine how function dimensionality affects the performance of our transfer learning approach. In 2-D (see Fig. 7 and Table I in the supplementary material [41]), the transferred model generally outperforms the model trained from scratch at small sample sizes (e.g., 50), with F23 and F24 as exceptions. This advantage mostly holds with more data, though F19 and F24 also become exceptions. The transferred model consistently outperforms the original model, except on F23. In 10-D (see supplementary material [41]), the transferred RFR shows a clear advantage at small sizes (e.g., 50 samples), but the gap narrows as more data (e.g., 500 samples) becomes available. Transfer is less effective on complex, multi-modal functions like F16 and F21–F24, where performance remains limited.
@@ V-B Transferring RFR on Real-world applications
Figure 4: The comparison evaluates Random forest regression models obtained through transfer learning against those trained from scratch on the Kinematics of a Robot Arm. It examines different transfer data sample sizes, with each figure cell showing the percentage difference in average SMAPE (%) between the two approaches for specific transfer settings. Positive values indicate better accuracy from transfer learning (i.e. lower SMAPE).
Figure 5: The comparison evaluates Random forest regression models obtained through transfer learning against those trained from scratch on the Real-World Optimization Benchmark in Vehicle Dynamics. It examines different transfer data sample sizes, with each figure cell showing the percentage difference in average SMAPE (%) between the two approaches for specific transfer settings. Positive values indicate better accuracy from transfer learning (i.e. lower SMAPE).
@@ V-B1 Porkchop Plot Benchmarks in Interplanetary Trajectory Optimization
Fig. 3 shows the SMAPE percentage differences between RFR models trained from scratch and those adapted via transfer learning across various transfer scenarios and dataset sizes. The transferred model typically outperforms the model trained from scratch when fewer than 100 samples are available, highlighting its advantage in data-scarce settings. As the transfer dataset grows, the performance gap narrows, with the models trained from scratch catching up. Fig. 10 (see supplementary material [41]) offers a detailed view of this trend. In most cases, the transferred model also outperforms the original RFR.
The original RFR model rarely surpasses the transferred or scratch-trained models when the transfer dataset is extremely limited, such as with just 10 samples, though occasional exceptions are observed. In such exceptions, the transferred model may overfit the limited data during optimization, leading to poor generalization to the broader target domain. Overfitting can be identified when the transferred model performs better than the original RFR on the transfer samples, as it is explicitly optimized for them. These findings highlight that while the transfer learning approach is generally robust, its effectiveness can be challenged in scenarios with severely restricted data availability. Similar performance patterns are observed in other experimental scenarios, such as Earth-to-Venus and Mercury-to-Earth transfers, as shown in Fig. 11, 12, 13, and 14, which are included in the supplementary material [41].